Find the hydrostatic pressure exerted on an object that is 4.45 m below the surface of gasoline and its absolute pressure. The density of gasoline is 680 kg/m3.
1
Expert's answer
2020-03-27T10:55:29-0400

Hydrostatic pressure equal: "p_h=\rho_ggh=680*9.81*4.45=30018.6" Pa

Absolute pressure equal: "p=p_0+p_h=101300+30018.6=131318.6" Pa